OpenStudy (anonymous):

One for Stephen Hawking! Thornes park is an existing development.A further site of 2.73 hectares is being offered. calculate the area of each plot using algebra. Plot A is twice as large as plot B. Plot B is 1.2 hectares smaller than plot C. The site is to be divided into 3 plots.

OpenStudy (anonymous):

Let one of the plot become x and form equations for each of three statements you were given.

OpenStudy (anonymous):

one of the plots*

OpenStudy (anonymous):

of the three*

OpenStudy (anonymous):

Let the size of Plot B be x. I will do the first statement for you and we will see if you can do the second one. Your first statement says that Plot A is twice the size of Plot B. And you know we just let Plot B equal x. So the statement can be reworded such as this: "Plot A is twice x." \[PlotA=2x\]

OpenStudy (anonymous):

What can you say about Plot C in terms of x. Try reading the 2nd statement over and over again until you get the gist of it.

OpenStudy (anonymous):

Plot C is 1.2 hectares bigger than B?

OpenStudy (anonymous):

C=B-1.2?

OpenStudy (anonymous):

If i let B be X c=x-1.2?

OpenStudy (anonymous):

When do you ever use this stuff in the real world...aggghhhh

OpenStudy (anonymous):

c=X + 1.2 I think ive got it?

OpenStudy (anonymous):

B = A divide by 2?

OpenStudy (anonymous):

or C - 1.2?

OpenStudy (anonymous):

A=2X B=A divide by 2 C=X + 1.2

OpenStudy (anonymous):

next step . (Xx2)+(x/2)+(x+1.2) ????

OpenStudy (anonymous):

Taken me ages . B + (B+1.2)+(2B)=2.73 4B+1.2=2.73 B=(2.73-1.2)/4 B=0.3825 A=0.765 C=1.5825 A+B+C= 2.73